If the intensity of radiation emitted by a surface covered with lamp black (α = 0.96) in the normal direction is 1.85 x 103

W/m2

Sr. Calculate the temperature of the surface if it follows Lambert’s Cosine Law


1
Expert's answer
2022-02-06T12:31:42-0500

Lambert’s cosine law states that the radiant intensity from the ideal diffusely reflecting surface and cosine of the angle θ between the direction of incident light and surface normal are directly proportional.

Heat radiated per second:


"E= \sigma AT^4"


"1.85\times10^3= 5.7\times10^-8 \times 1\times T^4"


"T =424.44 ^0K"